OpenStudy (anonymous):

MEDALLLLLLLLLL!!!!!!!!! The transformation from f to g represents a __________ stretch. f as a function of x is equal to the square root of x and g as a function of x is equal to the square root of quantity 0.1 times x

OpenStudy (dayakar):

\[f(x) = \sqrt{x}\] \[g(x) = 0.1\times \sqrt{x}\]

OpenStudy (anonymous):

is that a vertical stretch

OpenStudy (anonymous):

no its a vertical compression right?

OpenStudy (mathmale):

@hamza123: Your input, please? If in doubt, graph both f and g on the same set of coordinate axes.